Conducting Rails Two conducting rails 31.5 cm apart rest on a 4.97

Conducting Rails

Two conducting rails 31.5 cm apart rest on a 4.97
Answer #1

gravitational force = magnetic force

==> m g sin(theta) = B L i

==> m g sin(theta) = B L (emf/R)

==> m g sin(theta) = B L (B L v cos(theta)/R)

==> m g sin(theta) = B^2 L^2 (v cos(theta)/R)

==> v = (m g R sin(theta))/(B^2 L^2 cos(theta))

==> v = (0.0450*9.8*0.750*sin(4.97))/(0.700*0.700*0.315*0.315*cos4.97)

==> v = 0.592 m/s
Answer #2

formula for terminal velocity is

V = Rmg sin theta/(B^2L^2 cos^ theta)

where R is resisatnce L is Length B is mag field

so here

V = 0.75 * 0.045 * 9.8* sin 4.9/(0.7*0.315* cos 4.97)^2

V = 0.593 m/s